# DynEx: Dynamic Code Synthesis with Structured Design Exploration for Accelerated Exploratory Programming

### v1

@@ 3.2.1. Project Requirements
DynEx further enhances functionality with pre-approved libraries like Chart.js, an open source javascript library used for creating interactive and visually appealing charts on websites, and GoJs, as javascript library for creating interactive diagrams such as flow charts, mind maps, and process diagrams. A common issue with LLM code-generation tools is their difficulty in handling external libraries and packages. OpenHands can get stuck in an loop installing unusable packages, while tools like Claude Artifact fail when trying to import non-existent component libraries. These limitations pose significant challenges for generating reliable code with LLMs. To prevent this, we restrict our system to use only the MUI React Library and third-party API calls outside of OpenAI. However, this restriction can limit the complexity of generated applications, so we allow the use of Chart.js and GoJS for creating charts and diagrams. When our system recommends these libraries as requirements, we use few-shot code examples to ensure smooth integration. We can expand the list of pre-approved libraries this way to ensure safe and reliable code generation. For our example, neither Chart.js or GoJS is identified as a project requirement.
Project Specification
DynEx generates a project specification (spec) (see Fig  3- D), a detailed overview of the prototype idea which connects the concepts from the Design Matrix to flesh out the application layout, user interaction, and logic of the app. The spec transforms abstract ideas into actionable implementation instructions, ensuring the functionality of the app is well-aligned with the design; it provides a clear roadmap for development. Although there is potential overlap between the spec and Grounding level in the Design Matrix, the project specification focuses more on the technical requirements for the project. If placeholder data is needed, the spec helps define the schema. If GPT must be self-invoked, it details what the code should prompt for and the expected response. Ultimately, the spec drives the implementation plan; DynEx breaks the spec down into steps to iteratively implement on later on in the workflow.
In this example, the system determines that placeholder data and images are necessary components for the UI to be realized. The spec details that the system must generate placeholder data with fields that include Chinese words and phrases, images, and IDs. It also details that images must be returned for each quiz question to assist with visual learning.
Synthetic Placeholder Data
DynEx also enables the generation of synthetic placeholder data (see Fig  3- E). By using the spec and the Design Matrix as contextual inputs, it creates realistic placeholder data to serve as the foundation for building data-driven prototypes. We leverage few-shot examples with LLMs to guide the generation of relevant and appropriate placeholder data. The user can regenerate, approve, or modify the outputs. In our example, the user generates placeholder data, providing the prototype with a solid base of Chinese phrases to practice with, making this step essential for effective app development.
3.2.2. Step-by-step Implementation
Figure 4. DynEx User Process: The system first decomposes the project into steps that are modifiable by the user. The user iteratively implements each step. The user can add, remove, or update steps at any point. The backend LLM then generates the UI code – the user can also interact with the UI . The user can directly modify the code output, debug, and revert to previous versions. The user iterates through steps until the final prototype is complete
@@ 3.2.2. Step-by-step Implementation
A diagram of DynEx’s user process. The system first decomposes the project into steps that modifiable by the user. The user iteratively implements each step. The user can add, remove, or update steps at any point. The back-end LLM then generates the UI code – the user can also interact with the UI . The user can directly modify the code output, debug, and revert to previous versions. The user iterates through steps until the final prototype is complete
To implement the prototype, we employ a dynamic, iterative approach to development that allows for human direction at every step, which can be seen in Fig  4. The system breaks down the spec into steps, where each step is the next-smallest testable iteration of the previous step. In our example, the step list is broken down to:
  1. (1)
    ”Set up the React application and create the main layout with the three sections: ’Flashcards’, ’Quiz’, and ’Progress’”
  2. (2)
    ”Implement the ’Flashcards’ section”
  3. (3)
    ”Implement the ’Quiz’ section, utilizing the spaced repetition algorithm”
  4. (4)
    ”Implement the ’Progress’ section”
  5. (5)
    ”Call GPT to generate images for the Chinese characters”
The user can regenerate, approve, or modify this task list.
The system next executes steps sequentially. In our example, the system implements (1) ”Create the HTML structure”. The system generates the initial code, and the UI is rendered for the user to interact with. If there are failures, the user can debug by iteratively prompting the system in the right direction to regenerate code (see Fig  4 - I), redoing the step (see Fig  4 - H), or debugging the code by hand (see Fig  4 - K). Once the user confirms that it works, they can proceed to (2) ”Implement the ‘Flashcards section’.” The user does not modify the step and clicks the button to generate the code. Once the user tests and approves step 2, they can move on to (3) ”Implement the ‘Quiz’ section.”
Debugging and Iteration within Steps
At (3) ”Implement the ‘Quiz’ section,” there’s an error with the “submit” button in the ‘Quiz’ section. Instead of displaying the next quiz question, it displays both the next quiz question and the next quiz question’s answer. To fix this, the user can make use of the “Iterate” box to debug the program by explaining the bug through the prompt (see Fig  4 - I). The prompts the system to fix this bug: “After I click submit, it moves on to the next question but also shows the answer of the next question,” and clicks “Iterate”. After generating the code, the debugged UI and code is shown on the system. Users can iterate as many times as they want and switch between versions until they are happy with the final step output(see Fig  4 - J).
Feature Adding
The user tests and approves step 3, and moves on to steps (4) ”Implement the ‘Progress’ section” and (5) ”Call GPT to generate images.” Thus, the user has completed the original implementation plan. Dynex enables users to add more features as they test prototypes. After using the app and testing it out, the user decides they would like another feature: in the flashcards section where they practice the words, they would like a visual indicator to show which words they have mastered, and which words they are not familiar with. The user adds a step 7 to do so (see Fig  4 - G).
Version Control
By breaking our project into distinct steps, where users can update, remove, and add features, we enable version control. Each step is saved, creating a clear version history that captures the most advanced working prototype while offering the flexibility to iterate between versions. This means that users can easily revert to prior working states if they want to reconsider a feature, without worrying if leftover changes will affect the process.
Since LLM code-generation can sometimes produce faulty or unusable code, this stepwise approach serves as a safety net where users can always fall back on the last successful version, ensuring that they always have a stable point to return to. While this is not a main contribution of our system, it’s essential to support version control in exploratory programming to provide structure and safety in an unpredictable process, protect users from irreversible errors, and support smooth iteration.
3.3. Creating Multiple Prototypes
Once a prototype is created, users can create additional prototypes by revisiting the Design Matrix, tweaking parameters, exploring variations, and comparing the results (see Fig  4 - B). The different prototypes are displayed in the sidebar of DynEx (see Fig  4 - A), allowing users to easily revisit any prototype, pick up where they left off, and continue iterating. Users can modify entries in the Design Matrix and re-implement new designs to generate a diverse set of prototypes. This is a consistent method with exploratory programming, empowering users to explore multiple solutions in parallel and refine their designs through hands-on iteration.
4. Evaluation
@@ 4.1.1. Participants
Our evaluation was conducted through a qualitative study of 10 programmers (6 male, 4 female); there were 6 industry software developers and 4 CS students, all with many years of technical experience. They were recruited through a snowball sampling of local university students and alumni. Information regarding the participants age, role, and years of technical exposure are included in Table  1. Participants were paid \$20 per hour for their time, and the study was conducted with each participant for approximately 90 minutes.
@@ 4.1.3. Claude Artifact
We compared DynEx against Claude Artifact, an industry standard tool, as the baseline. Claude Artifact is a chatbot, where every prompt elicits a UI and non-editable code. The chat screen is on the left, and the Artifact to interact with the UI is on the right. The user can prompt the chatbot to brainstorm more ideas or specify more features – an example can be seen in Figure  5.
Because Claude Artifact can generate functioning code from short prompts, it is a very powerful tool for prototyping. Claude Artifact is also widely accessible and creates visually-appealing UIs. Although it is limited to 4096 output tokens (approximately 450 lines of code), Claude Artifact is a strong baseline because it still produces rich prototypes and stores long chat histories that track how the user further developed their application from the initial idea.
We considered other tools as the baseline, including OpenHands and GPTPilot. We found OpenHands to be less reliable than Claude Artifact in producing functional prototypes. In our exploration with GPTPilot, we found its user interactions limited to technical specifications. It did not allow for sufficient design space exploration, which is largely emphasized in DynEx. Since the user has the ability to engage with design space exploration through the flexibility of the chatbot interface, we decided Claude Artifact was the most appropriate baseline for our study.
4.2. Design Exploration through the Design Matrix
[…]
@@ 4.2.1. Procedure
An experimenter first explained the concepts of DynEx through a slide deck that introduced design exploration and exploratory programming, focusing on user-centric design principles and LLM code-generation. The experimenter then gave a demonstration of both systems. For Claude Artifact, the experimenter demonstrated how to prompt it to create a UI and add features through the chatbot. For DynEx, the experimenter demonstrated how to use the Design Matrix and iteratively implement, debug, and add features to their prototype. After the demonstrations for each system, participants were given 30 minutes to create a prototype using both DynEx and Claude Artifact. Half of the participants started prototyping with Claude Artifact, and the other half started with DynEx. After the conclusion of the two tasks, an experimenter conducted a semi-structured interview to understand participant experiences.
@@ 4.3.2. RQ2: Convergence - To what extent does DynEx allow users to better develop their idea?
Participants indicated that they were able to sufficiently build off of their own ideas using DynEx. P9 stated that DynEx allowed them to ”incorporate [their] initial idea and helped [them] refine it.” Participants also reported that using DynEx’s Design Matrix allowed them to better specify the problem that they were trying to solve, specifically noting that the Person dimension was helpful to inform the remaining entries of the matrix. Participants also felt that they were better able to explore the solution space with DynEx as compared to Claude Artifact. Overall, DynEx was as very helpful during the design process; P8 stated that they ”didn’t think many [questions were] left unanswered.” From these results, we conclude that DynEx is more successful at developing users ideas.
@@ 4.3.3. RQ3: Implementation - To what extent does DynEx enable the code to realize a complex idea?
Divergent Thinking: To what extent were you able to explore different conceptual options using system X? 4.2 6.1 0.0066 -3.52
@@
5x5 table displaying average score given by participants for either system for four questions (one title row) measuring the system’s ability to support exploratory programming. Each question was rated on a scale of 1 (Bad) to 7 (Good). The four questions are 1. Divergent Thinking: To what extent were you able to explore different conceptual options using system X?. 2. Convergent Thinking: To what degree were you able to better develop/specify your idea using system X? 3. Idea Realization: To what degree was the code able to realize your idea when using system X? 4. Application Complexity: How complex is the prototype produced by system X? Columns are: the question, average score for Claude Artifact, average score for DynEx, p-value from paired t-test on difference of system scores, t-statistic from the same test. The differences for the first, second, and fourth questions are statistically significant. The difference for the third question is not.
@@ 4.3.4. RQ4: Overall - To what extent does DynEx allow for a better prototyping experience?
Overall, we found that DynEx was more successful for exploratory programming compared to Clauade Artifact. In terms of success at accomplishing the task at hand (NASA-TLX performance), participants on average scored DynEx 5.9/7 compared to only 4.2/7 for Claude Artifact (see Table 3). 8/10 participants rated DynEx a 6/7 or higher in terms of NASA-TLX performance. This difference in rating was statistically significant at the p = 0.05 level (see Table  3).
The difference in ratings between DynEx and Claude Artifact for the remaining NASA-TLX metrics (mental demand, physical demand, temporal demand, effort, and frustration) were statistically insignificant (see Table  3). However, there were still interesting findings for the mental effort and demand metrics. The NASA-TLX score for both had a large range of responses; many participants found that DynEx gave them freedom to think conceptually, but also required them to read and wait more as compared to Claude Artifact. However, while using Claude Artifact, users expressed that it required more mental exertion to brainstorm ideas and guide the prototypes development direction. Hence, depending upon user’s preferences (i.e., do they prefer reading LLM-generations or writing prompts themselves), scores for mental exertion may vary. This same discrepancy likely explains the lack of statistical significance in the difference in ratings for the NASA-TLX effort category. From these results, we conclude that DynEx more successfully allowed users to create a final application as compared to the Claude Artifact baseline.
@@ 4.4.1. DynEx’s Design Matrix inspired new solutions
DynEx’s Design Matrix enabled participants to consider ideas they had not thought of, broadening their initial concept and allowing them to adequately explore the design space. P3 prototyped an app that recommended restaurants based on friends reviews. She noted: ”It was really cool that it gave me ideas – otherwise I would have gone with what I had imagined …[initially, which was not as good]”. She had pictured a typical, feed-like-layout, but through DynEx’s exploration, ended up prototyping a card-swipe interaction that she enjoyed more.
P5 experienced something similar, stating that ”For every [entry in the matrix], it included [ideas] that I wanted, but also considered options that I hadn’t considered but wanted to build…it did a good job of articulating ideas that I had in my mind, touching on tangential pieces that I might be interested in but didn’t know how to bring together.” P5 created a addiction recovery application, and was particularly intrigued by using cognitive behavioral theory (CBT) principles brainstormed in the Approach dimension. He ultimately created an app that allowed him to journal triggers, relapses, and good days, utilizing CBT help identify unhelpful thoughts.
P8 created a prototype to aggregate concert tickets based on value-for-money – he had been thinking about this idea for the past two years, but the Design Matrix was able to help him make progress with new insights. He stated,”even though I kind of knew what I wanted to do, I really wouldn’t have thought of these steps so fast. Stuff like: ’do you want a centralized database?’… [for even] simple things like that, I would have to think for a much longer time if I wanted to build it myself… it saved me from spending a lot of time on unimportant things… and gave me conceptual freedom.”.
On the contrary, Claude Artifact did not facilitate the exploration of the design space, but rather, implemented prototypes exactly as the user prompted. P2 noted that ”with DynEx, there’s a lot more time put into the [end-user of the prototype] and [a lot more] feature-oriented thoughts…[with Claude], I didn’t know what to [prompt] at times… since there were no recommendations.” Additionally, P4 stated that when using Claude Artifact, ”[he] was just thinking on [his] own… it did not come up with any of the ideas, [he] came up with all the ideas”. The Design Matrix gave users a better starting foundation; P9 stated that DynEx allowed him to ”incorporate [his] initial idea and refine it… the brainstorming helped [him] think through rough points [he] had missed out on”. Exploring the design space prevented users from design fixation, allowed them to ideate on new solutions, and challenged their existing thinking.
@@ 4.4.2. DynEx’s prototypes are more feature-rich and intuitive
Users were also pleasantly surprised by DynEx’s ability to realize more intuitive and realistic applications. DynEx was able to generate more relevant placeholder data and emulate a true user experience. P2 noted that ”having the faked data really helps with understanding what is going on with the system and interacting with it”, compared to Claude Artifact’s ”very simple” prototype with limited placeholder data. P3, who created the restaurant recommendation app, stated that she could not emulate a user experience from her prototype with Claude Artifact because ”[her] data could not mimic it… [it didn’t even] have reviews”; in comparison, DynEx’s placeholder data ”helped a lot [for her to envision] the UI properly”.
@@
P8 also noted the intuitive features in his concert aggregator prototype. With DynEx, the UI was formatted like a table, with important information surfaced – the main user interaction prioritized searching for valuable concert tickets with low cognitive load. In contrast, the UI for Claude Artifact had a feed view, with lots of information about each concert and ticket, which P8 disliked after interacting with the prototype. P8 greatly preferred DynEx’s prototype, stating: ”It was intuitive. It was feature-rich. It had all the important features, like sorting by columns and sorting by genre, dates, value-for-money etc…I liked the conciseness of the information… contrary to [Claude Artifact’s] prototype which was… not-fit for this use case.”. Ultimately, users were satisfied with the cohesion between features in DynEx-created prototypes.

### v3

@@ 5.2.1. Project Scoping
DynEx further enhances prototype functionality with pre-approved libraries. DynEx’s code outputs all use Material UI (MUI), a comprehensive React component library offering foundational elements like buttons, dropdowns, input forms, and more to create visually appealing applications. Additionally, DynEx’s project requirements allow the selection of Chart.js, an open source Javascript library used for creating interactive charts on websites, and GoJs, as Javascript library for creating interactive diagrams such as flow charts, mind maps, and process diagrams. When the system recommends these libraries as requirements, we use few-shot code examples to ensure smooth integration. The list of pre-approved libraries can be expanded this way to ensure safe and reliable code generation.
5.2.2. Modular Stepwise Implementation
DynEx employs few-shot examples to generate the plan, which can be found in Appendix D. The few-shot examples direct the system to generate a plan that mirrors how a developer would naturally approach the task – where each step is modular and progressively adds features to the prototype, building on what has already been implemented. The plan has approximately 3-6 steps, depending on the complexity of the application. At each step, DynEx calls Claude’s API to generate code. To maintain continuity, the prompt explicitly instructs Claude to retain code from previous steps. DynEx also employs a function to clean up any natural language in the code response, ensuring the output is executable. The code rules used in the prompt can be found in Appendix E.
Debugging and Iteration within Steps.
LLM code-generation is not perfect, and there are many ways that DynEx supports error recovery. The user has the option to regenerate the code within each step, manually edit the generated code and preview its output, or rollback to a previous step with working code. DynEx also offers an “Iterate” box to debug the program by explaining the bug through the prompt. The system will reissue a query to Claude, prompting it to fix the buggy snippet of code while keeping the remaining code intact. The prompt can be seen in Appendix H.
5.2.3. Self-invoked Multi-Modal LLMs
Dynex self-invokes GPT-4 and DALL-E 2 to enable dynamic and flexible content creation based on the project requirements. If GPT or Images is selected as a requirement, we use few-shot code examples to guide the self-invocation. In the few-shot prompt, we provide multiple Javascript code snippets that use a try-catch block to make asynchronous API calls to GPT-4 and DALL-E 2. Each example includes appropriate parameters, such as the OpenAI API endpoint URL, headers for authorization (including the API key), and necessary model identifiers. These snippets guide the generation of code that properly demonstrate how to structure safe OpenAI API calls. Examples can be found in Appendix F and G.
6. Evaluation
@@ 6.1.1. Participants
Our evaluation was conducted through a qualitative study of 10 programmers (6 male, 4 female); there were 6 industry software developers and 4 CS students. They were recruited through a snowball sampling of local university students and alumni. Information regarding the participants age, role, and years of technical exposure are included in Table  1. Participants were paid \$20 per hour for their time, and the study was conducted with each participant for approximately 90 minutes.
@@ 6.2.2. Claude Artifact
We compared DynEx against Claude Artifact, an industry standard tool, as the baseline. Claude Artifact is a chatbot, where every prompt elicits a UI and non-editable code. The chat screen is on the left, and the Artifact to interact with the UI is on the right. The user can prompt the chatbot to brainstorm more ideas or specify more features – an example can be seen in Figure  7.
We considered other tools as the baseline, including OpenHands and GPTPilot. We found OpenHands to be less reliable than Claude Artifact in producing functional prototypes. In our exploration with GPTPilot, we found its user interactions limited to technical specifications. It did not allow for sufficient design space exploration, which is largely emphasized in DynEx.
We found Claude Artifact to be the best available baseline for our evaluation. Because Claude Artifact can generate functioning code from short prompts, it is a very powerful tool for prototyping. Claude Artifact is also widely accessible and creates visually-appealing UIs. Although it is limited to 4096 output tokens (approximately 450 lines of code), it still produces rich prototypes and stores long chat histories that track how the user further developed their application from the initial idea. While it lacks an explicitly structured framework for design exploration, Claude Artifact has all the capabilities for people to explore the dimensions of a design space before generating their code. Its organic structure empowers users to iteratively refine their designs, and identify and address gaps in their UI prototypes as they emerge. This reactive exploration mirrors the natural way people interact with LLMs and offers an intuitive way to discover new possibilities within a design space.
@@ 6.2.3. Procedure
An experimenter first explained the concepts of DynEx through a slide deck that introduced design exploration and exploratory programming, focusing on user-centric design principles and LLM code-generation. The experimenter then gave a demonstration of both systems. For Claude Artifact, the experimenter demonstrated how to prompt it to create a UI and add features through the chatbot. For DynEx, the experimenter demonstrated how to use the Design Matrix and iteratively implement, debug, and add features to their prototype. After the demonstrations for each system, participants were given 30 minutes to create a prototype using DynEx and 30 minutes to create a prototype using Claude Artifact. Half of the participants started prototyping with Claude Artifact, and the other half started with DynEx. Participants were not allowed to modify the spec, because the spec is a natural translation of the Design Matrix into a technical implementation description. After the conclusion of the two tasks, an experimenter conducted a semi-structured interview to understand participant experiences.
@@ 6.3.2. RQ2: Convergence - To what extent does DynEx allow users to better develop their idea?
Participants indicated that they were able to sufficiently build off of their own ideas using DynEx. P9 stated that DynEx allowed them to ”incorporate [their] initial idea and helped [them] refine it.” Participants also reported that using DynEx’s Design Matrix allowed them to better specify the problem that they were trying to solve, specifically noting that the Person dimension was helpful to inform the remaining entries of the matrix. Participants also felt that they were better able to explore the solution space with DynEx as compared to Claude Artifact. Overall, DynEx was very helpful during the design process; P8 stated that they ”didn’t think many [questions were] left unanswered.” From these results, we conclude that DynEx is more successful at developing users ideas.
@@ 6.3.3. RQ3: Implementation - To what extent does DynEx enable the code to realize a complex idea?
Divergent Thinking: To what extent were you able to explore different conceptual ideas? using system X? 4.2 6.1 0.0066 -3.52
@@
5x5 table displaying average score given by participants for either system for four questions (one title row) measuring the system’s ability to support exploratory programming. Each question was rated on a scale of 1 (Bad) to 7 (Good). The four questions are 1. Divergent Thinking: To what extent were you able to explore different conceptual ideas? using system X?. 2. Convergent Thinking: To what degree were you able to better develop/specify your idea using system X? 3. Idea Realization: To what degree was the code able to realize your idea when using system X? 4. Application Complexity: How complex is the prototype produced by system X? Columns are: the question, average score for Claude Artifact, average score for DynEx, p-value from paired t-test on difference of system scores, t-statistic from the same test. The differences for the first, second, and fourth questions are statistically significant. The difference for the third question is not.
@@ 6.3.4. RQ4: Overall - To what extent does DynEx allow for a better prototyping experience?
Overall, we found that DynEx was more successful for exploratory programming compared to Clauade Artifact. In terms of success in accomplishing the task at hand (NASA-TLX performance), participants on average scored DynEx 5.9/7 compared to only 4.2/7 for Claude Artifact (see Table 3). 8/10 participants rated DynEx a 6/7 or higher in terms of NASA-TLX performance. This difference in rating was statistically significant at the p = 0.05 level (see Table  3).
The difference in ratings between DynEx and Claude Artifact for the remaining NASA-TLX metrics (mental demand, physical demand, temporal demand, effort, and frustration) were statistically insignificant (see Table  3). However, there were still interesting findings for the mental effort and demand metrics. These NASA-TLX scores had a large range of responses; many participants found that DynEx gave them freedom to think conceptually, but also required them to read and wait longer compared to Claude Artifact. However, users expressed that Claude Artifact required more mental exertion to brainstorm ideas and guide the prototype’s development direction. Hence, depending upon user’s preferences (i.e., do they prefer reading LLM-generations or writing prompts themselves), scores for mental exertion may vary. This same discrepancy can explain the lack of statistical significance in the difference in ratings for the NASA-TLX effort category. From these results, we conclude that DynEx more successfully allowed users to create a final application as compared to the Claude Artifact baseline.
@@ 6.4.1. DynEx’s Design Matrix inspired new solutions
DynEx’s Design Matrix enabled participants to consider ideas they had not thought of, broadening their initial concept and allowing them to adequately explore the design space. P3 prototyped an app that recommended restaurants based on friends reviews. She noted: ”It was really cool that it gave me ideas — otherwise I would have gone with what I had imagined … [initially, which was not as good]”. She had pictured a typical, feed-like-layout, but through DynEx’s exploration, ended up prototyping a card-swipe interaction that she enjoyed more.
P5 experienced something similar, stating that ”For every [entry in the matrix], it included [ideas] that I wanted, but also considered ideas that I hadn’t considered but wanted to build…it did a good job of articulating ideas that I had in my mind, touching on tangential pieces that I might be interested in but didn’t know how to bring together.” P5 created a addiction recovery application, and was particularly intrigued by using cognitive behavioral theory (CBT) principles brainstormed in the Approach dimension. He ultimately created an app that allowed him to journal triggers, relapses, and good days, utilizing CBT help identify unhelpful thoughts.
P8 created a prototype to aggregate concert tickets based on value-for-money — he had been thinking about this idea for the past two years, but the Design Matrix was able to help him make progress with new insights. He stated,”even though I kind of knew what I wanted to do, I really wouldn’t have thought of these steps so fast. Stuff like: ’do you want a centralized database?’… [for even] simple things like that, I would have to think for a much longer time if I wanted to build it myself… it saved me from spending a lot of time on unimportant things… and gave me conceptual freedom.”.
On the contrary, Claude Artifact did not facilitate the exploration of the design space, but rather, implemented prototypes exactly as the user prompted. P2 noted that ”with DynEx, there’s a lot more time put into the [end-user of the prototype] and [a lot more] feature-oriented thoughts…[with Claude], I didn’t know what to [prompt] at times… since there were no recommendations.” Additionally, P4 stated that when using Claude Artifact, ”I was just thinking on my own… it did not come up with any of the ideas, I came up with all the ideas”. The Design Matrix gave users a better starting foundation; P9 stated that ”[DynEx helped me] incorporate my initial idea and refine it… the brainstorming helped me think through rough points [he] had missed out on”. Exploring the design space prevented users from design fixation, allowed them to ideate on new solutions, and challenged their existing thinking.
@@ 6.4.2. DynEx’s prototypes are more feature-rich and intuitive
Users were also pleasantly surprised by DynEx’s ability to realize more intuitive and realistic applications. DynEx was able to generate more relevant placeholder data and emulate a true user experience. P2 noted that ”having the faked data really helps with understanding what is going on with the system and interacting with it”, compared to Claude Artifact’s ”very simple” prototype with limited placeholder data. P3, who created the restaurant recommendation app, noted that she could not emulate a user experience from her prototype with Claude Artifact, stating ”its data could not mimic it… [it didn’t even] have reviews”; in comparison, DynEx’s placeholder data ”helped a lot [to envision] the UI properly”.
@@
P8 also noted the intuitive features in his concert aggregator prototype. With DynEx, the UI was formatted like a table, with important information surfaced — the main user interaction prioritized searching for valuable concert tickets with low cognitive load. In contrast, the UI for Claude Artifact had a feed view, with lots of information about each concert and ticket, which P8 disliked after interacting with the prototype. P8 greatly preferred DynEx’s prototype, stating: ”It was intuitive. It was feature-rich. It had all the important features, like sorting by columns and sorting by genre, dates, value-for-money etc…I liked the conciseness of the information… contrary to [Claude Artifact’s] prototype which was… not-fit for this use case.”. Ultimately, users were satisfied with the cohesion between features in DynEx-created prototypes.